OpenStudy (anonymous):

Jack usually mows his lawn in 6 hours. Marilyn can mow the same yard in 5 hours. How much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

\[\frac{6\times 5}{6+5}\]

OpenStudy (amistre64):

if jack can eat a whole pie in 6 minutes; and Marilyn can eat a whole pie in 5 minutes; how much of the pie will the eat together in 1 minute?

OpenStudy (anonymous):

mmmm pie

OpenStudy (amistre64):

pecan coconut lemon marange

OpenStudy (amistre64):

the problems boil down to finding how much work each can do in 1 hour (min, sec, week, etc) and then adding them together to find out the total that can be done in 1 hour

OpenStudy (anonymous):

I got 30/11, but that can't be the right answer. It must be 2 hours and something...

OpenStudy (amistre64):

.... if only there was a way to convert 30/11 into some sort of mixed system of numbers lol

OpenStudy (anonymous):

it is \[\frac{30}{11}\] now convert to a mixed number. 11 goes in to 30 twice, with a remainder of 8, so the mixed number is \[2\tfrac{8}{11}\]

OpenStudy (anonymous):

ahhhh

OpenStudy (anonymous):

Thanks!